OpenStudy (anonymous):

what is the slope and y-intercept of 18x+4y=112

OpenStudy (anonymous):

Rearrange the equation according to y = mx + c .

OpenStudy (anonymous):

y = 28 - 4.5x

OpenStudy (anonymous):

m= -9/2 b=28